OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME.... 2|z+1|-3=z-2 solve

OpenStudy (asnaseer):

do you know what the "|" symbols around an expression mean?

OpenStudy (anonymous):

absolute value

OpenStudy (asnaseer):

correct, so this means what ever appears between these symbols (positive or negative value), we just drop the sign of the result and just consider its magnitude. agreed?

OpenStudy (anonymous):

agreed

OpenStudy (asnaseer):

ok, so we can break these types of problems into two: 1. consider the case when the expression between the "|" symbols is positive 2. consider the case when the expression between the "|" symbols is negative

OpenStudy (asnaseer):

so lets take the 1st case, in this case (since the expression between the symbols is positive), the "|" can just be removed to give: 2(z+1) - 3 = z - 2 can you solve this to find z?

OpenStudy (anonymous):

2z +2-3=z-2 2z-1=z-2 z=-1

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

now lets look at the second case...

OpenStudy (asnaseer):

if the expression between the "|" symbols is negative, then taking its absolute value is the same as putting a minus sign in fron of it. e.g. |-3| = -(-3) = 3 understand?

OpenStudy (anonymous):

yup

OpenStudy (asnaseer):

great! so that gives us: 2(-(z+1)) - 3 = z - 2 can you solve this to find z?

OpenStudy (anonymous):

-2z+2-3=z-2 -2z-1=z-2 -1=3z-2 3z=1 z=1/3

OpenStudy (asnaseer):

almost - you made a sign error in the first step: 2(-(z+1)) - 3 = z - 2 2(-z-1) - 3 = z - 2 -2z -2 -3 = z - 2

OpenStudy (anonymous):

-2z -2 -3 = z - 2 -2z-5=z-2 -5=3z-2 -3=3z z=-1

OpenStudy (asnaseer):

perfect! so you see there is only ONE solution to both cases: z=-1 hope you understood the process. :)

OpenStudy (anonymous):

thank you very much!

OpenStudy (asnaseer):

yw :)